12. [1pt]
A piano wire is stretched to a tension of 330 N. A wire is just a string made of metal. If the wire is 1.76 m long and has a total mass of 0.250 kg, what is the phase speed of transverse waves travelling along this wire?
Answer #1

Linear mass density is

u = m/L =0.25/1.76

u =0.142 kg/m

Speed of wave is given by

V=\sqrt{\frac{T}{\mu }}=\sqrt{\frac{330}{0.142}}=48.2 m/s